### v1

# Deep Learning assisted Port-Cycling based Channel Sounding for Precoder Estimation in Massive MIMO Arrays

Abstract
Future wireless systems are expected to employ a substantially larger number of transmit ports for channel state information (CSI) estimation compared to current specifications. Although scaling ports improves spectral efficiency, it also increases the resource overhead to transmit reference signals across the time-frequency grid, ultimately reducing achievable data throughput. In this work, we propose an deep learning (DL)-based CSI reconstruction framework that serves as an enabler for reliable CSI acquisition in future 6G systems. The proposed solution involves designing a port-cycling mechanism that sequentially sounds different portions of CSI ports across time, thereby lowering the overhead while preserving channel observability. The proposed CSI Adaptive Network (CsiAdaNet) model exploits the resulting sparse measurements and captures both spatial and temporal correlations to accurately reconstruct the full-port CSI. The simulation results show that our method achieves overhead reduction while maintaining high CSI reconstruction accuracy.
I Introduction
The massive multiple-input multiple-output (MIMO) system is a key technology for wireless communication systems, offering benefits including spatial multiplexing, capacity gain, etc. To achieve these benefits, accurate CSI is vital, which incurs substantial reference signal overhead with increasing channel dimensions. In current systems, downlink channel acquisition is performed using CSI-RS [1]. It is useful in assessing channel conditions and reports back the channel quality, the number of independent streams, and the beamforming vector to be utilized [2]; however, its overhead grows linearly with the array size thereby reducing the resources available for data transmission.
To address the growing overhead, industry and academic efforts have explored Artificial Intelligence/ Machine Learning (AI/ML) techniques for CSI feedback. In particular, 3GPP has identified AI/ML based CSI compression and prediction as initial use-cases for AI/ML based air interface [3]. From an academic perspective, CsiNet was the pioneer in demonstrating effective non-linear compression of CSI with reduced payload sizes [4]. In [5], authors have investigated CSI extrapolation across time, frequency and spatial domains to actively work on feedback reduction. While these approaches reduce feedback overhead or solve temporal issues, they do not address the instantaneous overhead associated with sounding a large number of antenna ports, which remains a main bottleneck in massive MIMO systems.
Motivated by this observation, this work focuses on reducing instantaneous channel sounding overhead in the antenna dimension, while maintaining compatibility with current standard compliant transmission. Unlike the environment based channel prediction approach that rely on site-specific information (i.e. images) [6], we adopt a port-cycling based channel sounding strategy in which a subset of ports are activated at a particular time instant, and the full panel is sequentially covered over multiple cycles.
Building on this framework, we propose CsiAdaNet, a deep learning based CSI reconstruction model that exploits the temporal and spatial correlations across the temporally aggregated CSI measurements. By coherently aggregating these sparse observations across time, CsiAdaNet predicts the quantized Type-II precoder defined in 5G NR [1],[7],enabling precoder estimation without requiring full-port instantaneous CSI.
The paper is organized as follows: The system model, port cycling scheme are introduced in Section-II. In Section-III, the details behind the CsiAdaNet model are described. The simulation setup and the results are described in Section-IV. The conclusions are summarized in Section-V.
II System Model
Consider a downlink massive MIMO system with $N_{t}\gg 1$ transmit antennas with a uniform planar array at the base station (BS) and $N_{r}$ receive antennas at the UE. The system is operated in orthogonal frequency division multiplexing (OFDM) with $K$ subcarriers. The received signal at the $k^{th}$ subcarrier can be expressed as
$\mathbf{y}_{k}=\mathbf{h}_{k}\mathbf{v}^{H}_{k}\mathbf{x}_{k}+\mathbf{n}_{k},$ (1)
where $\mathbf{y}_{k}\in\mathbb{C}^{N_{r}\times 1}$ denotes the received signal, $\mathbf{h}_{k}\in\mathbb{C}^{N_{r}\times N_{t}}$ denotes the channel matrix, $\mathbf{v}_{k}\in\mathbb{C}^{N_{L}\times N_{t}}$ denotes the precoding vector, $\mathbf{x}_{k}\in\mathbb{C}^{N_{L}\times 1}$ denotes the data vector and $n_{k}$ denotes the additive noise of the $k^{th}$ subcarrier. The BS is equipped with $N_{t}=2\times N_{x}\times N_{y}$ antennas, where $N_{x},N_{y}$ denotes the number of dual polarized antenna elements at the horizontal and vertical dimensions. The channel matrix $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times K}$ is obtained by stacking the channel matrices of each CSI subcarrier respectively.
II-A Type II codebook in 5G NR
5G-NR defines the configurable Type-II of codebooks based on a 2D-DFT based grid of beams in horizontal and vertical dimensions of the antenna array. It selects the optimal $L$ orthogonal beams and linearly combines them in such a way that it maximizes the energy in the eigen vector direction. The Rel.15 Type-II codebook [1] for a given layer $\nu$ and the entire wideband is defined as:
$\mathbf{W}^{(\nu)}=\boldsymbol{\Lambda}^{(\nu)}\left[\begin{array}[]{c}\displaystyle\sum_{i=0}^{L-1}{v}_{l^{(i)}m^{(i)}}p_{\nu,i}\varphi_{\nu,i}\\[10.0pt] \displaystyle\sum_{i=0}^{L-1}{v}_{l^{(i)}m^{(i)}}p_{\nu,i+L}\varphi_{\nu,i+L}\end{array}\right]$ (2)
Quantity Description
$v_{l,m}$ IDFT beam, $v_{l,m}=v_{l}\otimes u_{m}$
$v_{l}$: horizontal beam; $u_{m}$: vertical beam
$L$ Number of beams, $L\in\{2,3,4\}$
$p_{\nu}$ Wideband amplitude
$\varphi_{\nu}$ Wideband phase
$\Lambda^{(\nu)}$ Normalization factor
$\Lambda^{(\nu)}=\frac{1}{\sqrt{N_{x}N_{y}\sum_{l=0}^{2L-1}(p_{l})^{2}}}$
TABLE I: Rel.15 Type-II quantities
  • Beam Set: It refers to the over sampled beam group out of the available $O_{1}O_{2}$ groups.
  • Beam Indices: It refers to the top-$L$ orthogonal beams out of $N_{1}N_{2}$ beams within the selected beamset.
  • Wideband Amplitude: It corresponds to reporting the quantized wideband amplitudes for the selected 2L12L-1 orthogonal beams. The strongest beam within the beam set is normalized to unit amplitude. The quantized amplitudes are drawn from the set, $P={\{0,\sqrt{\frac{1}{64}},\sqrt{\frac{1}{32}},\sqrt{\frac{1}{16}},\sqrt{\frac{1}{8}},\sqrt{\frac{1}{4}},\sqrt{\frac{1}{2}},1\}}$
  • Wideband Phase: This corresponds to the reporting of wideband phase of the selected 2L12L-1 orthogonal beams. The strongest beam within the beam set is normalized to zero phase. The quantized phase alphabets are drawn from the set $\Phi=\{e^{\frac{j2\pi n}{N_{psk}}},n\in\{0,1,...,N_{psk}-1\}\}$
The beams are selected from within a beamset based on Algorithm 1 [8] and the wideband components are computed based on Algorithm 2.
Algorithm 1 Beam set and Beam Indices Selection
1:Input: Over-sampled beam grid $\{\mathbf{V}_{b}\}_{b=1}^{B}$, covariance $\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$, $L$ : Number of beams
2:Output: Beam Set ($B^{\star}$), Beam Indices ($\mathcal{I}^{\star}$)
3:
$\mathbf{R}=\begin{bmatrix}\mathbf{R}_{11}&\mathbf{R}_{12}\\ \mathbf{R}_{21}&\mathbf{R}_{22}\end{bmatrix}$
4:for $b=1$ to $B$ do
5:  $\mathbf{r}\leftarrow\mathbf{V}_{b}^{H}(\mathbf{R}_{11}+\mathbf{R}_{22})\mathbf{V}_{b}$
6:  $\mathbf{s}\leftarrow|\operatorname{diag}(\mathbf{r})|$
7:  $\mathcal{I}_{b}\leftarrow\operatorname{TopL}(\mathbf{s})$
8:  $\eta_{b}\leftarrow\sum_{i\in\mathcal{I}_{b}}s_{i}$
9:end for
[…]
11:return $B^{\star},\mathcal{I}^{\star}$
Algorithm 2 Wideband Amplitude and Phase Selection
1:Input: Beamset $b^{\star}$, beam indices $\mathcal{I}^{\star}$, covariance matrix $\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$
2:Output: Wideband Amplitude ($\mathcal{P}_{\nu}^{\star}$), Wideband Phase ($\mathcal{\varphi}_{\nu}^{\star}$)
[…]
5:for $l=1$ to $L$ do
6:  $p^{\nu}_{0,l}\leftarrow\mathbf{e}_{0,\nu}^{H}\mathbf{v}_{l}$, $p^{\nu}_{1,l}\leftarrow\mathbf{e}_{1,\nu}^{H}\mathbf{v}_{l}$
7:end for
8:for $\text{pol}\in\{0,1\}$ do
9:  $a^{\nu}_{pol}\leftarrow|p^{\nu}_{pol}|$
10:  $k^{\star}_{\nu,pol}\leftarrow\arg\max_{k}a^{\nu}_{k,pol}$
11:  $\tilde{{a}^{\nu}_{pol}}\leftarrow{a}^{\nu}_{pol}/a_{k^{\star}_{\nu,pol}}$
12:  $\mathbf{A}_{\nu,pol}\leftarrow|\tilde{\mathbf{a}^{\nu}_{pol}}|,\quad\boldsymbol{\theta}_{\nu,pol}\leftarrow\angle\tilde{\mathbf{a}^{\nu}_{pol}}$
13:  for $l=1$ to $L$ do
14:   $\mathcal{A}^{\nu}_{pol,l}\leftarrow\arg\min_{i}\left|P_{i}-{A}^{\nu}_{pol,l}\right|$
15:   ${\Theta}^{\nu}_{pol,l}\leftarrow\arg\min_{j}\left|\Phi_{j}-{\theta}^{\nu}_{pol,l}\right|$
16:  end for
17:end for
[…]
19:return $\mathcal{P}_{\nu}^{\star}$, $\mathcal{\varphi}_{\nu}^{\star}$
Algorithm 3 Sub-panel Partitioning
1:Input: UPA size $(N_{x},N_{y})$, Number of partitions $\rho=(\rho_{x}\times\rho_{y})$
2:Output: Set of sub-arrays $\mathcal{S}$
[…]
5:for $i=0$ to $\rho_{x}-1$ do
6:  for $j=0$ to $\rho_{y}-1$ do
7:   $\mathcal{S}_{i,j}\leftarrow$ antenna elements in rows $iN_{x}^{\prime}:(i{+}1)N_{x}^{\prime})]$ and cols $jN_{y}^{\prime}:(j{+}1)N_{y}^{\prime})]$
8:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{\mathcal{S}_{i,j}\}$
9:  end for
10:end for
11:return $\mathcal{S}$
II-B Sub-array partitioning
The main motivation in this work is to utilize a lesser number of ports for active transmission of reference signals. To reduce the number of active ports for transmission, the full-planar array is partitioned into smaller sub-arrays. In general, the active subset of ports can be selected using sub-panel, interleaved or arbitrary configurations. In this work, we opt to model the sub-array based on sub-panel partitioning, as it preserves the antenna spacing, thereby avoiding spatial aliasing in the beam computation.
The uniform planar array of size ($N_{x}\times N_{y}$) is divided into $\rho=\rho_{x}\times\rho_{y}$ contiguous sub-panels, each consisting of $N^{\prime}_{x}=N_{x}/\rho_{x}$ and $N^{\prime}_{y}=N_{y}/\rho_{y}$ antenna elements. This results in $\rho$ non-overlapping sub-arrays, each containing $N^{\prime}_{t}=N^{\prime}_{x}\times N^{\prime}_{y}\times N_{pol}$ ports as mentioned in Algorithm 3. and an example configuration is illustrated in Fig. 1.
Figure 1: Illustration of sub-panel based partitioning
II-C Port Cycling Scheme
It is a method of cycling through the available subset of ports at different time instants for downlink reference signal transmission. The motivation is that, when beam related quantities such as beamset ($\mathcal{B}$), dominant beam index ($\mathcal{I}_{0}$) do not vary across each of the sub-array along time, it underlines the fact that wireless channel statistics such as dominant path, angular spread remain relatively stable within the transmission duration. A score is defined to measure the degree of variation of each of the beam related metric ($\mathcal{M}$).
$\text{Variation Score}=\frac{1}{\rho-1}\sum_{i=1}^{\rho-1}\left|\mathcal{M}^{\,i+1}-\mathcal{M}^{\,i}\right|.$
Figure 2: Plot of variation scores
As shown in Fig. 2, both the beamset index and dominant beam index exhibit low variation across port-cycling instances, particularly at moderate-to-high SNR region. The beamset remains highly stable even under noisy conditions, while the beam index converges quickly with increasing SNR. These observations validate the feasibility of temporally aggregating port-cycled CSI measurements to predict precoder.
Figure 3: Model Architecture of CSIAdaNet
III CsiAdaNet
This section provides details about a Deep Learning assisted solution for the port cycling based system. The proposed CSIAdaNet is an DL model that resides at the user equipment (UE) side and operates on $\rho$ periodic CSI measurements, each corresponding to a distinct sub-panel based sub-array with a periodicity of $P$ ms. After cycling through all the $\rho$ sub-panels and estimating the channel corresponding to them, the model predicts the quantized Type-II codebook parameters.
At a given prediction instance, the operation of CsiAdaNet can be expressed as
$B^{\star},\;\mathcal{I}^{\star},\;\mathcal{P}^{\star},\mathcal{\varphi}^{\star}=\mathcal{F}(\{\mathcal{H}_{t+iP}\}_{i=0}^{\rho-1};\theta)$ (3)
where $\mathcal{F}$ denotes the DL model, $\theta$ refers to the model parameters and $\mathcal{H}_{t+iP}$ denoting the CSI corresponding to sub-array sounded at time $t+iP$.
The model performs the predictions sequentially with contextual dependency. The beamset is initially predicted, followed by the prediction of the top $L$ beam indices conditioned on the beamset. The predicted beamset and beam indices are then jointly used to estimate the quantized wideband amplitude levels and phase levels, enabling a hierarchical passage of information that remains consistent with the practical codebook selection strategies.
The input to the model consists of $\rho$ sub-array CSI tensors each represented as $\mathcal{H}_{t}\in\mathbb{R}^{N_{x}^{\prime}\times N_{y}^{\prime}\times(N_{r}N_{sc})\times(2N_{pol})}$, where $N_{x}^{\prime}$$N_{y}^{\prime}$ denote the number of horizontal and vertical antenna elements in the sub-array, $N_{r}$ is the number of receive antennas, $N_{sc}$ is the number of subcarriers configured for CSI in the entire bandwidth, and $N_{pol}$ = 2 denotes the dual polarization. The factor of $2$ corresponds to the real and imaginary components of the complex channel co-efficients.
Each CSI tensor is independently processed by a shared encoder module ($\mathrm{Enc}$) to extract the spatial-frequency features. The input is arranged to explicitly preserve the sub-array structure $(N^{\prime}_{x}\times N^{\prime}_{y})$, enabling the convolutional layers to effectively learn channel variations across the sub-panel dimensions. The encoder consists of three consecutive 3D convolutional layers with filter sizes {16,32,64} followed by a Global Average Pooling operation to capture the effective features. It is finally passed through a Dense layer of $D=32$ units, yielding a compact latent representation
$\mathbf{z}_{t}=\mathrm{Enc}(\mathcal{H}_{t}),\;\mathbf{z}_{t}\in\mathbb{R}^{D}$ (4)
To exploit the temporal structure captured by the sequential sub-array sounding, the latent vectors $\{\mathbf{z}_{t}\}_{t=1}^{\rho}$ are passed through Gated Recurrent Unit (GRU).
The model architecture of CsiAdaNet is illustrated in the Fig. 3, employs multiple heads each corresponding to a distinct CSI feedback quantity. All of the heads operate on a shared latent representation obtained after temporal aggregation from the GRU layers and feature refinement through Head Block units. Each prediction is implemented using a lightweight feed-forward network comprising of a Dense layer followed by a softmax activation, thereby formulating the outputs as a classification over the respective discrete levels.
Beam Set Head: It predicts the over-sampled beamset index ($\mathcal{B}$), which is a coarse representation of the spatial direction and rotation in the beam space. The head operates on the shared representation and outputs a probability distribution over the $O_{1}\times O_{2}$ beamset levels.
Beam Indices Head: It predicts the top $L$ beam indices ($\mathcal{I}$) among the $N_{x}\times N_{y}$ orthogonal beams within the selected beamset. It takes in the beamset information in its latent form as an additional information to predict the beam indices accurately.
Wideband Amplitude Head: It predicts the quantized wideband amplitude levels associated with the selected top $L$ beams in the previous stage. The amplitudes capture the power distribution among the selected $2L$ beams and are drawn from a pre-defined finite set $P$.
Wideband Phase: It predicts the quantized wideband phase levels corresponding to the selected beams. The phase capture a way to coherently combine the selected beams to focus energy in the desired direction. The latent beam information from beamset and beam indices prediction are used in predicting the quantized phase levels.
The CSIAdaNet model is trained using a weighted sum of Sparse-Categorical Cross-Entropy (SCCE) losses for individual CSI quantities. For a single sample, the loss is defined as
[…]
where $\hat{y}$ denotes the ground-truth class index and $\mathit{p}_{x}$ is the predicted class probability score. The task specific loss is given by
$\mathcal{L}_{x}=\mathbb{E}[\mathit{l}_{SCCE}(\hat{x},\mathit{p}_{x})],\;x\in\{\mathcal{B},\mathcal{I},\mathcal{P},\mathcal{\varphi}\}$ (6)
The overall training loss is defined as
$\mathcal{L}_{total}=\sum_{x\in\{\mathcal{B},\mathcal{I},\mathcal{P},\mathcal{\varphi}\}}\lambda_{x}\mathcal{L}_{x}$ (7)
The predicted codebook can be reconstructed as
$\mathbf{\tilde{W}}^{(\nu)}=\boldsymbol{\Lambda}^{(\nu)}\left[\begin{array}[]{c}\displaystyle\sum_{i=0}^{L-1}{V_{b}[\mathcal{B}^{\star}][\mathcal{I}^{\star}_{i}]}\times\mathcal{P}^{\star}_{\nu,i}\times e^{j\mathcal{\varphi}^{\star}_{\nu,i}}\\[8.0pt] \displaystyle\sum_{i=0}^{L-1}V_{b}[\mathcal{B}^{\star}][\mathcal{I}^{\star}_{i}]\times\mathcal{P}^{\star}_{\nu,i+L}\times e^{j\mathcal{\varphi}^{\star}_{\nu,i+L}}\end{array}\right]$ (8)
IV Simulation and results
In this section, we discuss the simulation framework considered. We use a 3GPP standards compliant proprietary simulator based on 5G NR Rel.19.
IV-A Link-Level Simulation Assumptions
The simulator is configured to operate on 3GPP Urban Macro (UMa) Line-of-Sight (LoS) scenario with an inter-site distance of 200m [9]. It is a single cell setup, sectorized into three sectors. The BS is equipped with 128 antennas in a uniform planar array structure while the UE is configured with 4 antennas. The BS antenna array is partitioned into sub-panel based sub-arrays with spatial domain reduction factor ($\rho=4$). The periodicity ($P$) of downlink CSI-RS is set to $20ms$. In this simulation, we consider the rank of the channel be to one i.e., $\nu=1$. The key simulation parameters are summarized in Table II.
Parameter Value
Deployment Scenario Urban Macro (UMa) LoS
Number of transmit antennas ($N_{t}$) $128$
Transmit antenna array configuration $N_{x}=8,\;N_{y}=8,\;N_{pol}=2]$
Number of receive antennas ($N_{r}$) 44
CSI-RS periodicity ($T_{{csi}}$) $20\;\text{ms}$
Carrier frequency ($f_{c}$) $3.5\;\text{GHz}$
System bandwidth (BW) $10\;\text{MHz}$
Subcarrier spacing (scsscs) $15\;\text{kHz}$
UE velocity 30km/h30\;km/h
Spatial reduction factor ($\rho$) $4~(\rho_{x}=2,\;\rho_{y}=2)$
Number of antennas per sub-array ($N_{t}^{{}^{\prime}}$) 32
TABLE II: Simulation Assumptions
IV-B CsiAdaNet Model
The dataset consists of $30,000$ samples and is split into training, validation, and test sets in a $70:20:10$ ratio. The model is trained under multiple noise conditions by uniformly applying SNR levels from $\{-5,0,5,10,15\}$dB to the training data to increase its robustness. During testing, the model is evaluated independently at each SNR level in $\{-20,-15,-10,-5,0,5,10,15,20\}$ dB to assess performance across a wide SNR range. A batch size of $256$ is considered for $250$ epochs using the Adam optimizer with a learning rate of $10^{-3}$. The CsiAdaNet model is trained using a weighted multi-task loss, with a loss weights ($\lambda_{\mathcal{B}}:0.75,\lambda_{\mathcal{I}}:1.5,\;\lambda_{\mathcal{P}}:0.75,\;\lambda_{\mathcal{\varphi}}:1.25)$. A higher weightage is assigned to beam indices ($\lambda_{\mathcal{I}}$) to emphasize beam selection, as any error in that propagates directly in precoder construction. All of the prediction heads employ sparse categorical cross-entropy losses with sparse categorical accuracy as the evaluation metric. Early stopping is deployed with a patience of 25 and adaptive learning rate reduction are used to stabilize training and avoid overfitting.
To perform prediction, each sub-array is sounded at a different time instant, covering the full array over $\rho$ occasions. During training, all possible sounding orders are to be considered to avoid bais towards a specific sequence; For $\rho=4$ this results in $4!=24$ permutations. The port cycled estimates collected over $(\rho\times P)\;ms$, forms the input to the DL model, ensuring generalization against sounding order variations.
IV-C Metrics
As a traditional method, we consider the actual Type-II precoder obtained using the full set of ports ($W_{T2}$) as the reference for comparison, with the precoder quantities using Algorithms 1 and 2.
Squared Generalized Cosine Similarity(SGCS): It captures the directional alignment between the Type-II precoder traditional or CsiAdaNet predicted, and the dominant eigen vector of the corresponding covariance matrix with the actual number of ports of the channel. It ranges between [0,1].
$\mathrm{SGCS}(\mathbf{v}_{\text{eig}},\mathbf{w})=\frac{\left|\mathbf{v}_{\text{eig}}^{\mathrm{H}}\mathbf{w}\right|^{2}}{\lVert\mathbf{v}_{\text{eig}}\rVert_{2}^{2}\,\lVert\mathbf{w}\rVert_{2}^{2}},\mathbf{w}\in[\ \tilde{W},W_{T2}]\$(9)
Beamforming Gain (BF): It measures the effective power captured by the Type-II precoder traditional or CsiAdaNet based, relative to the optimal eigen-beamformer.
$Gain(\mathbf{v}_{\mathbf{w},\text{eig}})=\frac{\mathbf{w}^{H}\mathbf{R}\;\mathbf{w}}{\mathbf{v}_{\mathrm{eig}}^{H}\mathbf{R}\;\mathbf{v}_{\mathrm{eig}}\,},\mathbf{w}\in[\ \tilde{W},W_{T2}]\$(10)
(a) Cosine Similarity
(b) Beamforming Gain
(c) Beamset Accuracy
(d) Beam Indices Accuracy
(e) Wideband Amplitude Accuracy
(f) Wideband Phase Accuracy
Figure 4: Performance evaluation of CSI Type-II precoder reconstruction across SNR: (a) cosine similarity, (b) beamforming gain, (c) beamset accuracy, (d) beam indices accuracy, (e) wideband amplitude accuracy, and (f) wideband phase accuracy.
where $\mathbf{v}_{eig}$ denotes the dominant eigen vector, $\tilde{W}$ refers to the CsiAdaNet predicted precoder from port-cycled channel measurements and $W_{T2}$ refers to the actual Type-II precoder corresponding to the actual number of ports.
IV-D Performance Evaluation and Analysis
IV-D1 Precoder Alignment and Beamforming Performance
Figs. 5(a) and 5(b) evaluate the alignment and beamforming performance of the CsiAdaNet predicted precoder. As shown in Fig. 5(a), the DL based approach closely follows the traditional technique in terms of SGCS across all SNR regimes. At low SNR, the SGCS is lower (i.e., <0.2<0.2) due to noise-dominated sub-array CSI, but steadily improves with SNR and saturates at around $0.7$ in medium-to-high SNR regime, approaching the baseline trend. Consistently, Fig. 5(b) shows that beamforming gain achieved using CsiAdaNet predicted precoder closely matches the traditional baseline, with a small gap at high SNR (within $\approx 0.03$). The drop at low SNR arises from the higher sensitivity of beamforming gain to angular and amplitude estimation errors.
IV-D2 Analysis of CSI feedback quantities
Figs. 5(c)–5(f) summarize the prediction accuracy of the individual CSI feedback quantities estimated by CSIAdaNet. In Fig. 5(c), beamset prediction is highly reliable even at moderate SNR levels, which can be attributed to the fact that its a coarse spatial resolution and has inherent robustness to noise and spatial sparsity. This supports the assumption that dominant spatial characteristics remain stable over the port-cycling duration. In Fig. 5(d), the beam indices prediction improves more gradually with SNR, reflecting the increased sensitivity associated with finer angular resolution. The wideband amplitude accuracy in Fig. 5(e) follows a similar trend, as amplitude estimation depends on the correctness of the selected beam indices. Finally, Fig. 5(f) indicates that wideband phase prediction is the a challenging task, particularly at low SNR, due to its reliance on coherent combining. Nevertheless, accurate phase estimation is achieved at higher SNR levels, demonstrating the effectiveness of temporally aggregated port-cycled CSI measurements.
IV-D3 Complexity Analysis
This subsection compares the computational complexity of the proposed CSIAdaNet model with the conventional covariance based methods. Conventional approaches compute eigenvalue decomposition (EVD) of the full covariance matrix $R_{N_{t}}\in\mathbb{C}^{N_{t}\times N_{t}}$, incurring a complexity of $O(N^{3}_{t})$, along with an additional, $O(O_{1}O_{2}N^{2}_{t})$ operations for selecting beam related metrics. In contrast, the proposed DL based approach performs inference on sub-panel observations, that scales as a $O(\rho D^{2})$, where $\rho$ denotes the number of subpanels and $D$ is the sub-panel feature dimension. The proposed approach remains independent of the full array size $N_{t}$, providing improved scalability for large-array systems.
V Conclusion
This work presents a DL-assisted port-cycling based precoder estimation framework that jointly exploits spatial diversity across antennas and temporal diversity enabled by channel stationarity to reduce instantaneous CSI transmission overhead in large MIMO systems. Rather than sounding all ports simultaneously, spatially disjoint sub-panels are activated sequentially, and the resulting partial CSI measurements are coherently aggregated over time. The proposed CsiAdaNet model leverages the resulting spatial–temporal correlations to reconstruct Type-II codebook quantities from reduced CSI observations. Link-level simulations conducted in a 3GPP-compliant setup demonstrate that the proposed approach achieves near-optimal beam alignment and beamforming gain at moderate to high SNR levels, with only marginal degradation under noise-limited regions. Although all antenna ports are eventually sounded across multiple occasions, the number of active ports at any given instant is reduced by a factor of $\rho$, leading to reduced instantaneous overhead and RF chain utilization. This distinction between instantaneous and aggregate CSI overhead highlights a key design consideration for next-generation large-scale MIMO systems, and demonstrates the effectiveness of exploiting spatial and temporal diversity for scalable and energy-efficient CSI acquisition.

### v2

@@
 Future wireless systems are expected to employ a substantially larger number of transmit ports for channel state information (CSI) estimation compared to current specifications.
-Although scaling ports improves spectral efficiency, it also increases the resource overhead to transmit reference signals across the time-frequency grid, ultimately reducing achievable data throughput.
+Although this improves spectral efficiency, it also increases the resource overhead to transmit reference signals across the time-frequency grid, ultimately reducing achievable data throughput.
 In this work, we propose an deep learning (DL)-based CSI reconstruction framework that serves as an enabler for reliable CSI acquisition in future 6G systems.
@@
 I Introduction
-The massive multiple-input multiple-output (MIMO) system is a key technology for wireless communication systems, offering benefits including spatial multiplexing, capacity gain, etc.
-To achieve these benefits, accurate CSI is vital, which incurs substantial reference signal overhead with increasing channel dimensions.
-In current systems, downlink channel acquisition is performed using CSI-RS [1].
-It is useful in assessing channel conditions and reports back the channel quality, the number of independent streams, and the beamforming vector to be utilized [2]; however, its overhead grows linearly with the array size thereby reducing the resources available for data transmission.
+Massive multiple-input multiple-output (MIMO) system is a crucial technology for wireless communication systems, offering benefits including spatial multiplexing, capacity gain, etc.
+These benefits rely on accurate CSI, whose acquisition incurs substantial overhead with increasing channel dimensions.
+In systems operating in frequency division duplexing (FDD) scenario, downlink channel acquisition is performed using CSI-RS [1], which enables reporting of channel quality, number of independent streams, and the beamforming vector [2].
+However, CSI resource overhead grows linearly with the array size, reducing data throughput.
 To address the growing overhead, industry and academic efforts have explored Artificial Intelligence/ Machine Learning (AI/ML) techniques for CSI feedback.
-In particular, 3GPP has identified AI/ML based CSI compression and prediction as initial use-cases for AI/ML based air interface [3].
-From an academic perspective, CsiNet was the pioneer in demonstrating effective non-linear compression of CSI with reduced payload sizes [4].
-In [5], authors have investigated CSI extrapolation across time, frequency and spatial domains to actively work on feedback reduction.
-While these approaches reduce feedback overhead or solve temporal issues, they do not address the instantaneous overhead associated with sounding a large number of antenna ports, which remains a main bottleneck in massive MIMO systems.
+In particular, 3GPP has identified CSI compression and prediction as initial use-cases for AI/ML based air interface [3].
+In academia, CsiNet demonstrated effective non-linear CSI compression [4], while subsequent works investigate CSI extrapolation across time, frequency and spatial domains [5].
+However these approaches primarily reduce feedback or temporal overhead and do not address the instantaneous overhead present due to large scale antenna sounding.
 Motivated by this observation, this work focuses on reducing instantaneous channel sounding overhead in the antenna dimension, while maintaining compatibility with current standard compliant transmission.
@@
 images) [6], we adopt a port-cycling based channel sounding strategy in which a subset of ports are activated at a particular time instant, and the full panel is sequentially covered over multiple cycles.
-Building on this framework, we propose CsiAdaNet, a deep learning based CSI reconstruction model that exploits the temporal and spatial correlations across the temporally aggregated CSI measurements.
-By coherently aggregating these sparse observations across time, CsiAdaNet predicts the quantized Type-II precoder defined in 5G NR [1],[7],enabling precoder estimation without requiring full-port instantaneous CSI.
+Building on this framework, we propose CsiAdaNet, a deep learning model that exploits the temporal and spatial correlations across the temporally aggregated CSI measurements to predict the quantized Type-II precoder defined in 5G NR [1],[7],enabling precoder estimation without requiring full-port instantaneous CSI.
+The main contributions of this work are summarized as follows:
+  1. 1.
+    A port-cycling based CSI sounding technique that reduces the instantaneous spatial overhead in massive MIMO systems.
+  2. 2.
+    A deep learning model that directly predicts Type-II precoder quantities from temporally aggregated port-cycled measurements.
+  3. 3.
+    Link-level simulation in a 3GPP compliant setup demonstrate proposed approach achieves near optimal performance.
 The paper is organized as follows: The system model, port cycling scheme are introduced in Section-II.
@@
 $\mathbf{y}_{k}=\mathbf{h}_{k}\mathbf{v}^{H}_{k}\mathbf{x}_{k}+\mathbf{n}_{k},$ (1)
-where $\mathbf{y}_{k}\in\mathbb{C}^{N_{r}\times 1}$ denotes the received signal, $\mathbf{h}_{k}\in\mathbb{C}^{N_{r}\times N_{t}}$ denotes the channel matrix, $\mathbf{v}_{k}\in\mathbb{C}^{N_{L}\times N_{t}}$ denotes the precoding vector, $\mathbf{x}_{k}\in\mathbb{C}^{N_{L}\times 1}$ denotes the data vector and $n_{k}$ denotes the additive noise of the $k^{th}$ subcarrier.
+where $\mathbf{y}_{k}\in\mathbb{C}^{N_{r}\times 1},\mathbf{h}_{k}\in\mathbb{C}^{N_{r}\times N_{t}},\mathbf{v}_{k}\in\mathbb{C}^{N_{L}\times N_{t}},\mathbf{x}_{k}\in\mathbb{C}^{N_{L}\times 1},n_{k}$ denotes the received signal, channel matrix, precoding vector, data vector and the additive noise of the $k^{th}$ subcarrier.
 The BS is equipped with $N_{t}=2\times N_{x}\times N_{y}$ antennas, where $N_{x},N_{y}$ denotes the number of dual polarized antenna elements at the horizontal and vertical dimensions.
-The channel matrix $\mathcal{H}\in\mathbb{C}^{N_{r}\times N_{t}\times K}$ is obtained by stacking the channel matrices of each CSI subcarrier respectively.
+The wideband covariance matrix is computed as
+$\mathbf{R}=\frac{1}{K}\sum_{i=1}^{k}h^{H}_{i}h_{i},\quad\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$ (2)
 II-A Type II codebook in 5G NR
@@
   • Wideband Amplitude: It corresponds to reporting the quantized wideband amplitudes for the selected 2L12L-1 orthogonal beams.
-The strongest beam within the beam set is normalized to unit amplitude.
 The quantized amplitudes are drawn from the set, $P={\{0,\sqrt{\frac{1}{64}},\sqrt{\frac{1}{32}},\sqrt{\frac{1}{16}},\sqrt{\frac{1}{8}},\sqrt{\frac{1}{4}},\sqrt{\frac{1}{2}},1\}}$
@@
 Algorithm 1 Beam Set and Beam Indices Selection
-1:Input: Over-sampled beam grid $\{\mathbf{V}_{b}\}_{b=1}^{B}$, covariance $\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$, $L$ : Number of beams
-2:Output: Beam Set ($B^{\star}$), Beam Indices ($\mathcal{I}^{\star}$)
-3:
-$\mathbf{R}=\begin{bmatrix}\mathbf{R}_{11}&\mathbf{R}_{12}\\ \mathbf{R}_{21}&\mathbf{R}_{22}\end{bmatrix}$
-4:for $b=1$ to $B$ do
-5:  $\mathbf{r}\leftarrow\mathbf{V}_{b}^{H}(\mathbf{R}_{11}+\mathbf{R}_{22})\mathbf{V}_{b}$
-6:  $\mathbf{s}\leftarrow|\operatorname{diag}(\mathbf{r})|$
-7:  $\mathcal{I}_{b}\leftarrow\operatorname{TopL}(\mathbf{s})$
-8:  $\eta_{b}\leftarrow\sum_{i\in\mathcal{I}_{b}}s_{i}$
-9:end for
+1:Input: $\{\mathbf{V}_{b}\}_{b=1}^{B}$, $\mathbf{R}$, $\mathbf{L}$
+2:Output: $B^{\star},\mathcal{I}^{\star}$
+3:for $b=1$ to $B$ do
+4:  $\mathbf{r}\!\leftarrow\!\mathbf{V}_{b}^{H}(\mathbf{R}_{11}+\mathbf{R}_{22})\mathbf{V}_{b}$
+5:  $\mathcal{I}_{b}\!\leftarrow\!\operatorname{Top}L(|\operatorname{diag}({r})|)$
+6:  $\eta_{b}\!\leftarrow\!\sum_{i\in\mathcal{I}_{b}}|{r}_{i}|$
+7:end for
 […]
-11:return $B^{\star},\mathcal{I}^{\star}$
+9:return $B^{\star},\mathcal{I}^{\star}$
 Algorithm 2 Wideband Amplitude and Phase Selection
-1:Input: Beamset $b^{\star}$, beam indices $\mathcal{I}^{\star}$, covariance matrix $\mathbf{R}\in\mathbb{C}^{N_{t}\times N_{t}}$
-2:Output: Wideband Amplitude ($\mathcal{P}_{\nu}^{\star}$), Wideband Phase ($\mathcal{\varphi}_{\nu}^{\star}$)
+1:Input: $b^{\star},\mathcal{I}^{\star},\mathbf{R}$Output: $\mathcal{P}^{\star}_{\nu},{\varphi}^{\star}_{\nu}$
 […]
+3:for $l=1$ to $L$ do
+4:  $p^{\nu}_{0,l}\!=\!\mathbf{e}^{H}_{0,\nu}\mathbf{v}_{l},\ p^{\nu}_{1,l}\!=\!\mathbf{e}^{H}_{1,\nu}\mathbf{v}_{l}$
+5:end for
+6:for $\text{pol}\!\in\!\{0,1\}$ do
+7:  $\tilde{\mathbf{a}}^{\nu}_{pol}\!\leftarrow\!|p^{\nu}_{pol}|/\max_{k}|p^{\nu}_{k,pol}|$
+8:  ${\theta}_{\nu,pol}\!\leftarrow\!\angle\tilde{\mathbf{a}}^{\nu}_{pol}$
+9:  $\mathcal{A}^{\nu}_{pol}\!\leftarrow\!\arg\min_{i}|P_{i}-\tilde{\mathbf{a}}^{\nu}_{pol}|$
+10:  ${\Theta}^{\nu}_{pol}\!\leftarrow\!\arg\min_{j}|\Phi_{j}-{\theta}_{\nu,pol}|$
+11:end for
 […]
-5:for $l=1$ to $L$ do
-6:  $p^{\nu}_{0,l}\leftarrow\mathbf{e}_{0,\nu}^{H}\mathbf{v}_{l}$, $p^{\nu}_{1,l}\leftarrow\mathbf{e}_{1,\nu}^{H}\mathbf{v}_{l}$
-7:end for
-8:for $\text{pol}\in\{0,1\}$ do
-9:  $a^{\nu}_{pol}\leftarrow|p^{\nu}_{pol}|$
-10:  $k^{\star}_{\nu,pol}\leftarrow\arg\max_{k}a^{\nu}_{k,pol}$
-11:  $\tilde{{a}^{\nu}_{pol}}\leftarrow{a}^{\nu}_{pol}/a_{k^{\star}_{\nu,pol}}$
-12:  $\mathbf{A}_{\nu,pol}\leftarrow|\tilde{\mathbf{a}^{\nu}_{pol}}|,\quad\boldsymbol{\theta}_{\nu,pol}\leftarrow\angle\tilde{\mathbf{a}^{\nu}_{pol}}$
-13:  for $l=1$ to $L$ do
-14:   $\mathcal{A}^{\nu}_{pol,l}\leftarrow\arg\min_{i}\left|P_{i}-{A}^{\nu}_{pol,l}\right|$
-15:   ${\Theta}^{\nu}_{pol,l}\leftarrow\arg\min_{j}\left|\Phi_{j}-{\theta}^{\nu}_{pol,l}\right|$
-16:  end for
-17:end for
-[…]
-19:return $\mathcal{P}_{\nu}^{\star}$, $\mathcal{\varphi}_{\nu}^{\star}$
-Algorithm 3 Sub-panel Partitioning
-1:Input: UPA size $(N_{x},N_{y})$, Number of partitions $\rho=(\rho_{x}\times\rho_{y})$
-2:Output: Set of sub-arrays $\mathcal{S}$
-[…]
-[…]
-5:for $i=0$ to $\rho_{x}-1$ do
-6:  for $j=0$ to $\rho_{y}-1$ do
-7:   $\mathcal{S}_{i,j}\leftarrow$ antenna elements in rows $iN_{x}^{\prime}:(i{+}1)N_{x}^{\prime})]$ and cols $jN_{y}^{\prime}:(j{+}1)N_{y}^{\prime})]$
-8:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{\mathcal{S}_{i,j}\}$
-9:  end for
-10:end for
-11:return $\mathcal{S}$
+13:return $\mathcal{P}^{\star}_{\nu},{\varphi}^{\star}_{\nu}$
 II-B Sub-array partitioning
-The main motivation in this work is to utilize a lesser number of ports for active transmission of reference signals.
-To reduce the number of active ports for transmission, the full-planar array is partitioned into smaller sub-arrays.
-In general, the active subset of ports can be selected using sub-panel, interleaved or arbitrary configurations.
-In this work, we opt to model the sub-array based on sub-panel partitioning, as it preserves the antenna spacing, thereby avoiding spatial aliasing in the beam computation.
-The uniform planar array of size ($N_{x}\times N_{y}$) is divided into $\rho=\rho_{x}\times\rho_{y}$ contiguous sub-panels, each consisting of $N^{\prime}_{x}=N_{x}/\rho_{x}$ and $N^{\prime}_{y}=N_{y}/\rho_{y}$ antenna elements.
-This results in $\rho$ non-overlapping sub-arrays, each containing $N^{\prime}_{t}=N^{\prime}_{x}\times N^{\prime}_{y}\times N_{pol}$ ports as mentioned in Algorithm 3.
-and an example configuration is illustrated in Fig.
+To reduce the number of active ports for sounding, the uniform planar array (UPA) of size ($N_{x}\times N_{y}$) is partitioned into $\rho=\rho_{x}\times\rho_{y}$ non-overlapping sub-panels.
+Each sub-panel consists of $N^{\prime}_{x}=N_{x}/\rho_{x}$ and $N^{\prime}_{y}=N_{y}/\rho_{y}$, resulting in $N^{\prime}_{t}=N^{\prime}_{x}\times N^{\prime}_{y}\times N_{pol}$ ports per sub-array.
+Sub-panel based partitioning, preserves the original antenna spacing, and thereby avoids spatial aliasing in the beam computation, making it well suited for port-cycling based CSI acquisition.
+An illustrative example of the procedure is shown in Fig.
 1.
@@
 II-C Port Cycling Scheme
-It is a method of cycling through the available subset of ports at different time instants for downlink reference signal transmission.
-The motivation is that, when beam related quantities such as beamset ($\mathcal{B}$), dominant beam index ($\mathcal{I}_{0}$) do not vary across each of the sub-array along time, it underlines the fact that wireless channel statistics such as dominant path, angular spread remain relatively stable within the transmission duration.
-A score is defined to measure the degree of variation of each of the beam related metric ($\mathcal{M}$).
+It refers to the sequential activation of non-overlapping sub-arrays across multiple time instances for downlink CSI transmission.
+At each instant, only one sub-array is sounded and the full antenna array is covered after $\rho$ cycles.
+This enables instantaneous reduction in spatial overhead while retaining the observability of the full array across time.
+The feasibility of this method relies on the temporal coherence of beam related quantities during the port-cycling duration.
+To quantify validity of this scheme, a variation score ($\mathcal{M}$) is defined as
+Figure 2: Plot of variation scores
 $\text{Variation Score}=\frac{1}{\rho-1}\sum_{i=1}^{\rho-1}\left|\mathcal{M}^{\,i+1}-\mathcal{M}^{\,i}\right|.$
-Figure 2: Plot of variation scores
-As shown in Fig.
-2, both the beamset index and dominant beam index exhibit low variation across port-cycling instances, particularly at moderate-to-high SNR region.
-The beamset remains highly stable even under noisy conditions, while the beam index converges quickly with increasing SNR.
-These observations validate the feasibility of temporally aggregating port-cycled CSI measurements to predict precoder.
-Figure 3: Model Architecture of CSIAdaNet
+Fig.
+2, illustrates that both beamset and dominant beam index exhibit low variation across port-cycling instances, particularly at moderate-to-high SNR region, validating the feasibility of temporally aggregating port-cycled CSI measurements to predict precoder.
 III CsiAdaNet
-This section provides details about a Deep Learning assisted solution for the port cycling based system.
-The proposed CSIAdaNet is an DL model that resides at the user equipment (UE) side and operates on $\rho$ periodic CSI measurements, each corresponding to a distinct sub-panel based sub-array with a periodicity of $P$ ms.
-After cycling through all the $\rho$ sub-panels and estimating the channel corresponding to them, the model predicts the quantized Type-II codebook parameters.
+This section presents a deep learning assisted solution for port cycling based system, where the proposed CSIAdaNet resides at the user equipment (UE) side and operates on $\rho$ periodic CSI sub-panel measurements, to predict the quantized Type-II codebook parameters.
 At a given prediction instance, the operation of CsiAdaNet can be expressed as
-$B^{\star},\;\mathcal{I}^{\star},\;\mathcal{P}^{\star},\mathcal{\varphi}^{\star}=\mathcal{F}(\{\mathcal{H}_{t+iP}\}_{i=0}^{\rho-1};\theta)$ (3)
-where $\mathcal{F}$ denotes the DL model, $\theta$ refers to the model parameters and $\mathcal{H}_{t+iP}$ denoting the CSI corresponding to sub-array sounded at time $t+iP$.
-The model performs the predictions sequentially with contextual dependency.
-The beamset is initially predicted, followed by the prediction of the top $L$ beam indices conditioned on the beamset.
-The predicted beamset and beam indices are then jointly used to estimate the quantized wideband amplitude levels and phase levels, enabling a hierarchical passage of information that remains consistent with the practical codebook selection strategies.
-The input to the model consists of $\rho$ sub-array CSI tensors each represented as $\mathcal{H}_{t}\in\mathbb{R}^{N_{x}^{\prime}\times N_{y}^{\prime}\times(N_{r}N_{sc})\times(2N_{pol})}$, where $N_{x}^{\prime}$$N_{y}^{\prime}$ denote the number of horizontal and vertical antenna elements in the sub-array, $N_{r}$ is the number of receive antennas, $N_{sc}$ is the number of subcarriers configured for CSI in the entire bandwidth, and $N_{pol}$ = 2 denotes the dual polarization.
-The factor of $2$ corresponds to the real and imaginary components of the complex channel co-efficients.
-Each CSI tensor is independently processed by a shared encoder module ($\mathrm{Enc}$) to extract the spatial-frequency features.
-The input is arranged to explicitly preserve the sub-array structure $(N^{\prime}_{x}\times N^{\prime}_{y})$, enabling the convolutional layers to effectively learn channel variations across the sub-panel dimensions.
-The encoder consists of three consecutive 3D convolutional layers with filter sizes {16,32,64} followed by a Global Average Pooling operation to capture the effective features.
-It is finally passed through a Dense layer of $D=32$ units, yielding a compact latent representation
-$\mathbf{z}_{t}=\mathrm{Enc}(\mathcal{H}_{t}),\;\mathbf{z}_{t}\in\mathbb{R}^{D}$ (4)
-To exploit the temporal structure captured by the sequential sub-array sounding, the latent vectors $\{\mathbf{z}_{t}\}_{t=1}^{\rho}$ are passed through Gated Recurrent Unit (GRU).
-The model architecture of CsiAdaNet is illustrated in the Fig.
-3, employs multiple heads each corresponding to a distinct CSI feedback quantity.
-All of the heads operate on a shared latent representation obtained after temporal aggregation from the GRU layers and feature refinement through Head Block units.
-Each prediction is implemented using a lightweight feed-forward network comprising of a Dense layer followed by a softmax activation, thereby formulating the outputs as a classification over the respective discrete levels.
-Beam Set Head: It predicts the over-sampled beamset index ($\mathcal{B}$), which is a coarse representation of the spatial direction and rotation in the beam space.
+$\{B^{\star},\;\mathcal{I}^{\star},\;\mathcal{P}^{\star},\mathcal{\varphi}^{\star}\}=\mathcal{F}(\{\mathcal{H}_{t+iP}\}_{i=0}^{\rho-1};\theta)$ (4)
+where $\mathcal{F}(\cdot)$ denotes the DL model, with parameters $\theta$ and $\mathcal{H}_{t+iP}$ denotes the CSI measurements corresponding to sub-array sounded at time $t+iP$.
+The model performs the predictions hierarchically with contextual dependency.
+The beamset is first predicted, followed by the top $L$ beam indices conditioned on the beamset, and finally the quantized wideband amplitude and phase levels.
+The input to the model consists of $\rho$ sub-array CSI tensors, $\mathcal{H}_{t}\in\mathbb{R}^{N_{x}^{\prime}\times N_{y}^{\prime}\times(N_{r}N_{sc})\times(2N_{pol})}$, where the factor of $2$ corresponds to the real and imaginary components of the complex channel co-efficients.
+Each CSI tensor is processed by a shared encoder module $\mathrm{Enc}(\cdot)$ to extract the spatial-frequency features while preserving the sub-array structure $(N^{\prime}_{x}\times N^{\prime}_{y})$.
+It comprises of three 3D convolutional layers with filter sizes {16,32,64} followed by a global average pooling and a dense layer of $D=32$ units, yielding a latent representation $\mathbf{z}_{t}=\mathrm{Enc}(\mathcal{H}_{t})\in\mathbb{R}^{D}$.
+The sequence $\{\mathbf{z}_{t}\}_{t=1}^{\rho}$ is further processed by a Gated Recurrent Unit (GRU) to capture temporal correlations across port-cycling instances and further feature refinement through Head Block units.
+The aggregated representation is then fed to multiple task-specific heads, each implemented as a feed-forward classifier operating on the shared latent features.The overall model architecture of CsiAdaNet is illustrated in Fig.
+3.
+The Beam Set Head predicts the over-sampled beamset index ($\mathcal{B}$), providing a coarse spatial representation .
 The head operates on the shared representation and outputs a probability distribution over the $O_{1}\times O_{2}$ beamset levels.
-Beam Indices Head: It predicts the top $L$ beam indices ($\mathcal{I}$) among the $N_{x}\times N_{y}$ orthogonal beams within the selected beamset.
-It takes in the beamset information in its latent form as an additional information to predict the beam indices accurately.
-Wideband Amplitude Head: It predicts the quantized wideband amplitude levels associated with the selected top $L$ beams in the previous stage.
-The amplitudes capture the power distribution among the selected $2L$ beams and are drawn from a pre-defined finite set $P$.
-Wideband Phase: It predicts the quantized wideband phase levels corresponding to the selected beams.
-The phase capture a way to coherently combine the selected beams to focus energy in the desired direction.
-The latent beam information from beamset and beam indices prediction are used in predicting the quantized phase levels.
-The CSIAdaNet model is trained using a weighted sum of Sparse-Categorical Cross-Entropy (SCCE) losses for individual CSI quantities. For a single sample, the loss is defined as
-[…]
-where $\hat{y}$ denotes the ground-truth class index and $\mathit{p}_{x}$ is the predicted class probability score. The task specific loss is given by
-$\mathcal{L}_{x}=\mathbb{E}[\mathit{l}_{SCCE}(\hat{x},\mathit{p}_{x})],\;x\in\{\mathcal{B},\mathcal{I},\mathcal{P},\mathcal{\varphi}\}$ (6)
+Conditioned on this, the Beam Indices Head predicts the top $L$ beam indices ($\mathcal{I}$) within the selected beamset.
+The Wideband Amplitude Head predicts the quantized wideband amplitude levels for the selected $2L$ beams from a pre-defined set $P$, while the Wideband Phase Head predicts the quantized phase levels from $\Phi$ for coherent beam combination.
+The CSIAdaNet is trained using a weighted multi-task sparse categorical cross-Entropy (SCCE) loss.
+For each CSI quantity ($x\in\{\mathcal{B},\mathcal{I},\mathcal{P},\mathcal{\varphi}\}$), the task specific loss is defined as $\mathcal{L}_{x}=\mathbb{E}[\mathit{l}_{SCCE}(\hat{x},\mathit{p}_{x})]$, where $\hat{x}$ denotes the ground-truth class index and $\mathit{p}_{x}$ is the predicted class probability score.
 The overall training loss is defined as
 $\mathcal{L}_{total}=\sum_{x\in\{\mathcal{B},\mathcal{I},\mathcal{P},\mathcal{\varphi}\}}\lambda_{x}\mathcal{L}_{x}$ (5)
-The predicted codebook can be reconstructed as
+Finally, the predicted codebook can be reconstructed as
 $\mathbf{\tilde{W}}^{(\nu)}=\boldsymbol{\Lambda}^{(\nu)}\left[\begin{array}[]{c}\displaystyle\sum_{i=0}^{L-1}{V_{b}[\mathcal{B}^{\star}][\mathcal{I}^{\star}_{i}]}\times\mathcal{P}^{\star}_{\nu,i}\times e^{j\mathcal{\varphi}^{\star}_{\nu,i}}\\[8.0pt] \displaystyle\sum_{i=0}^{L-1}V_{b}[\mathcal{B}^{\star}][\mathcal{I}^{\star}_{i}]\times\mathcal{P}^{\star}_{\nu,i+L}\times e^{j\mathcal{\varphi}^{\star}_{\nu,i+L}}\end{array}\right]$ (6)
+Figure 3: Model Architecture of CSIAdaNet
 IV Simulation and results
@@
 IV-A Link-Level Simulation Assumptions
-The simulator is configured to operate on 3GPP Urban Macro (UMa) Line-of-Sight (LoS) scenario with an inter-site distance of 200m [9].
-It is a single cell setup, sectorized into three sectors.
-The BS is equipped with 128 antennas in a uniform planar array structure while the UE is configured with 4 antennas.
-The BS antenna array is partitioned into sub-panel based sub-arrays with spatial domain reduction factor ($\rho=4$).
-The periodicity ($P$) of downlink CSI-RS is set to $20ms$.
-In this simulation, we consider the rank of the channel be to one i.e., $\nu=1$.
+The simulation is conducted using a 3GPP compliant link-level simulator under an Urban Macro (UMa) line-of-sight scenario with an 200m inter-site distance [9].
+A single-cell, three-sector deployment is considered, with a 128-antenna uniform planar array at the BS and four antennas at the UE.
+The BS array is partitioned into sub-panels with spatial reduction factor ($\rho=4$) and the CSI-RS periodicity is set to 20 ms, and single layer transmission is assumed, i.e ($\nu=1$).
 The key simulation parameters are summarized in Table II.
@@
 IV-B CsiAdaNet Model
-The dataset consists of $30,000$ samples and is split into training, validation, and test sets in a $70:20:10$ ratio.
-The model is trained under multiple noise conditions by uniformly applying SNR levels from $\{-5,0,5,10,15\}$dB to the training data to increase its robustness.
-During testing, the model is evaluated independently at each SNR level in $\{-20,-15,-10,-5,0,5,10,15,20\}$ dB to assess performance across a wide SNR range.
-A batch size of $256$ is considered for $250$ epochs using the Adam optimizer with a learning rate of $10^{-3}$.
-The CsiAdaNet model is trained using a weighted multi-task loss, with a loss weights ($\lambda_{\mathcal{B}}:0.75,\lambda_{\mathcal{I}}:1.5,\;\lambda_{\mathcal{P}}:0.75,\;\lambda_{\mathcal{\varphi}}:1.25)$.
-A higher weightage is assigned to beam indices ($\lambda_{\mathcal{I}}$) to emphasize beam selection, as any error in that propagates directly in precoder construction.
-All of the prediction heads employ sparse categorical cross-entropy losses with sparse categorical accuracy as the evaluation metric.
-Early stopping is deployed with a patience of 25 and adaptive learning rate reduction are used to stabilize training and avoid overfitting.
-To perform prediction, each sub-array is sounded at a different time instant, covering the full array over $\rho$ occasions.
-During training, all possible sounding orders are to be considered to avoid bais towards a specific sequence; For $\rho=4$ this results in $4!=24$ permutations.
-The port cycled estimates collected over $(\rho\times P)\;ms$, forms the input to the DL model, ensuring generalization against sounding order variations.
+The dataset consists of $30,000$ samples, split into training, validation, and test sets in a $70:20:10$ ratio.
+To improve robustness, training is performed under multiple SNR conditions uniformly sampled from $\{-5,0,5,10,15\}$ dB, while evaluation is conducted over a wider SNR range of 20-20 to $20$ dB.
+The model is trained for $250$ epochs using Adam optimizer with a learning rate of $10^{-3}$ and a batch size of $256$ employing a weighted multi-task loss, with a loss weights ($\lambda_{\mathcal{B}}:0.75,\lambda_{\mathcal{I}}:1.5,\;\lambda_{\mathcal{P}}:0.75,\;\lambda_{\mathcal{\varphi}}:1.25)$.
+A higher weightage is assigned to beam indices ($\lambda_{\mathcal{I}}$) to emphasize any error in that propagates directly in precoder construction.
+All heads use sparse categorical cross-entropy with accuracy as the evaluation metric, while early stopping and adaptive learning-rate reduction are applied to prevent overfitting.
+Sub-arrays are sequentially sounded for port cycling, with all ($\rho!$) sounding permutations used during training to ensure model robustness.
 IV-C Metrics
 As a traditional method, we consider the actual Type-II precoder obtained using the full set of ports ($W_{T2}$) as the reference for comparison, with the precoder quantities using Algorithms 1 and 2.
-Squared Generalized Cosine Similarity(SGCS): It captures the directional alignment between the Type-II precoder traditional or CsiAdaNet predicted, and the dominant eigen vector of the corresponding covariance matrix with the actual number of ports of the channel.
-It ranges between [0,1].
+Squared Generalized Cosine Similarity(SGCS): measures the directional alignment between the traditional or CsiAdaNet predicted Type-II precoder and the dominant eigen vector of the channel, with values ranging from 0 to 1.
 $\mathrm{SGCS}(\mathbf{v}_{\text{eig}},\mathbf{w})=\frac{\left|\mathbf{v}_{\text{eig}}^{\mathrm{H}}\mathbf{w}\right|^{2}}{\lVert\mathbf{v}_{\text{eig}}\rVert_{2}^{2}\,\lVert\mathbf{w}\rVert_{2}^{2}},\mathbf{w}\in[\ \tilde{W},W_{T2}]\$(7)
-Beamforming Gain (BF): It measures the effective power captured by the Type-II precoder traditional or CsiAdaNet based, relative to the optimal eigen-beamformer.
+Beamforming Gain (BF): measures the effective power captured by the traditional or CsiAdaNet based Type-II precoder, relative to the optimal eigen-beamformer.
 $Gain(\mathbf{v}_{\mathbf{w},\text{eig}})=\frac{\mathbf{w}^{H}\mathbf{R}\;\mathbf{w}}{\mathbf{v}_{\mathrm{eig}}^{H}\mathbf{R}\;\mathbf{v}_{\mathrm{eig}}\,},\mathbf{w}\in[\ \tilde{W},W_{T2}]\$(8)
@@
 Figs.
-5(a) and 5(b) evaluate the alignment and beamforming performance of the CsiAdaNet predicted precoder.
+5(a) and 5(b) evaluate the alignment and beamforming performance of CsiAdaNet.
 As shown in Fig.
-5(a), the DL based approach closely follows the traditional technique in terms of SGCS across all SNR regimes.
-At low SNR, the SGCS is lower (i.e., <0.2<0.2) due to noise-dominated sub-array CSI, but steadily improves with SNR and saturates at around $0.7$ in medium-to-high SNR regime, approaching the baseline trend.
+5(a), the DL based approach closely follows the conventional baseline across all SNRs, with SGCS increasing from below $0.2$ at low SNR to approximately $0.7$ in the medium-to-high SNR region.
 Consistently, Fig.
-5(b) shows that beamforming gain achieved using CsiAdaNet predicted precoder closely matches the traditional baseline, with a small gap at high SNR (within $\approx 0.03$).
-The drop at low SNR arises from the higher sensitivity of beamforming gain to angular and amplitude estimation errors.
+5(b) shows that beamforming gain achieved using CsiAdaNet, exhibiting close matches to the traditional baseline, with a small gap at high SNR (within $\approx 0.03$).
+The drop at low SNR is primarily due to noise-sensitive estimation.
 IV-D2 Analysis of CSI feedback quantities
 Figs.
-5(c)–5(f) summarize the prediction accuracy of the individual CSI feedback quantities estimated by CSIAdaNet.
+5(c)–5(f) summarize the prediction accuracy of the individual CSI feedback quantities.
 In Fig.
-5(c), beamset prediction is highly reliable even at moderate SNR levels, which can be attributed to the fact that its a coarse spatial resolution and has inherent robustness to noise and spatial sparsity.
-This supports the assumption that dominant spatial characteristics remain stable over the port-cycling duration.
+5(c), beamset prediction remains highly reliable even at moderate SNR levels, due to its coarse spatial resolution and inherent robustness to noise, validating the port-cycling assumption.
 In Fig.
-5(d), the beam indices prediction improves more gradually with SNR, reflecting the increased sensitivity associated with finer angular resolution.
+5(d), the beam indices accuracy improves more gradually with SNR, reflecting the finer angular sensitivity.
 The wideband amplitude accuracy in Fig.
-5(e) follows a similar trend, as amplitude estimation depends on the correctness of the selected beam indices.
+5(e) follows a similar trend, as amplitude estimation depends on the correctness of the beam selection.
 Finally, Fig.
-5(f) indicates that wideband phase prediction is the a challenging task, particularly at low SNR, due to its reliance on coherent combining.
-Nevertheless, accurate phase estimation is achieved at higher SNR levels, demonstrating the effectiveness of temporally aggregated port-cycled CSI measurements.
+5(f) indicates that wideband phase prediction is more challenging at low SNR but achieves high accuracy at moderate-to-high SNR, demonstrating the effectiveness of temporally aggregated port-cycled CSI.
 IV-D3 Complexity Analysis
-This subsection compares the computational complexity of the proposed CSIAdaNet model with the conventional covariance based methods.
-Conventional approaches compute eigenvalue decomposition (EVD) of the full covariance matrix $R_{N_{t}}\in\mathbb{C}^{N_{t}\times N_{t}}$, incurring a complexity of $O(N^{3}_{t})$, along with an additional, $O(O_{1}O_{2}N^{2}_{t})$ operations for selecting beam related metrics.
-In contrast, the proposed DL based approach performs inference on sub-panel observations, that scales as a $O(\rho D^{2})$, where $\rho$ denotes the number of subpanels and $D$ is the sub-panel feature dimension.
-The proposed approach remains independent of the full array size $N_{t}$, providing improved scalability for large-array systems.
+This subsection compares the computational complexity of the CsiAdaNet model with conventional methods.
+Conventional approaches require eigenvalue decomposition (EVD) of the full covariance matrix $R\in\mathbb{C}^{N_{t}\times N_{t}}$, incurring $O(N^{3}_{t})$ complexity.
+In contrast, CsiAdaNet performs inference on sub-panel observations, with complexity $O(\rho D^{2})$, independent of the full array size $N_{t}$, providing improved scalability for large-array systems.
 V Conclusion
-This work presents a DL-assisted port-cycling based precoder estimation framework that jointly exploits spatial diversity across antennas and temporal diversity enabled by channel stationarity to reduce instantaneous CSI transmission overhead in large MIMO systems.
-Rather than sounding all ports simultaneously, spatially disjoint sub-panels are activated sequentially, and the resulting partial CSI measurements are coherently aggregated over time.
-The proposed CsiAdaNet model leverages the resulting spatial–temporal correlations to reconstruct Type-II codebook quantities from reduced CSI observations.
-Link-level simulations conducted in a 3GPP-compliant setup demonstrate that the proposed approach achieves near-optimal beam alignment and beamforming gain at moderate to high SNR levels, with only marginal degradation under noise-limited regions.
-Although all antenna ports are eventually sounded across multiple occasions, the number of active ports at any given instant is reduced by a factor of $\rho$, leading to reduced instantaneous overhead and RF chain utilization.
-This distinction between instantaneous and aggregate CSI overhead highlights a key design consideration for next-generation large-scale MIMO systems, and demonstrates the effectiveness of exploiting spatial and temporal diversity for scalable and energy-efficient CSI acquisition.
+This work presents a DL-assisted port-cycling based precoder estimation that exploits both spatial diversity across antennas and temporal diversity arising from port-cycled measurements to reduce instantaneous CSI overhead in large MIMO systems.
+Rather than sounding all antenna ports simultaneously, spatially disjoint sub-panels are activated sequentially, and the partial CSI measurements are temporally aggregated .
+The proposed CsiAdaNet model leverages these spatial–temporal correlations to reconstruct Type-II codebook quantities from reduced CSI observations.
+Link-level simulations in a 3GPP-compliant setup demonstrate near-optimal beam alignment and beamforming gain at moderate to high SNR levels, with only marginal degradation under noisy regions.
+Although all antenna ports are eventually sounded across multiple occasions, the number of active ports at any given instant is reduced by a factor of $\rho$, reducing instantaneous overhead and RF chain usage, highlighting a key design consideration for scalable and energy-efficient CSI acquisition in next-generation large-scale MIMO systems.